Assume that Elodea leaf cells are rectangular and
Answer #1

A typical Elodea leaf cell is 0.05 mm long .

On this basis we will calculate the width and depth of the Elodea leaf cell:

Depth of one cell = 25% of length of one cell = 25% 0f 0.05 =0.0125 mm

Width of one cell = 25% of length of one cell = 25% 0f 0.05 =0.0125 mm

Volume of one Elodea leaf cell= length x width x depth =0.05x 0.0125x0.0125=7.8 x 10-6 mm3

Converting mm3 to micrometre3 : 1 mm3 = 109 micrometre3

Volume of one Elodea leaf cell=7.8 x 10-6 mm3 =7.8 x 10-6 x109 micrometre3=7.8x 103 micrometre3

Number of Elodea leaf cell in cm3 =1.0 cm3/7.8x 103 x10-12   =0.128x109

1 micrometre3 =10-12 cm3

Hence the number of cells would be grater than 1.0